OpenStudy (anonymous):

The Johnson Farm is divided into 15 equal sections. The house, barn, and fruit stand are in 3 of the sections. The remaining sections are orchards of fruit trees and vegetable gardens. What fraction is equal to the part of the farm that is orchards of fruit trees and vegetable. i know the answer is 4/5 but sumbody explain how

OpenStudy (anonymous):

\[\frac{15-3}{15}=\frac{4}{5} \]

OpenStudy (anonymous):

thank you robtobey